# SynCoGen: Synthesizable 3D Molecule Generation via Joint Reaction and Coordinate Modeling

### v1

@@ 4.2 Noising Schemes
Output: $\tilde{C}_{0}$ (re-centered ground truth), $C_{t}$ (interpolated noisy coords)
1:$\mathcal{D}_{t}\leftarrow\{\,i\mid X_{t}[i]\neq\pi_{X}\}$triangleright denoised blocks
2:$S_{t}[i,a]\leftarrow\mathbf{1}[i\notin\mathcal{D}_{t}\lor a\in\mathcal{A}_{i}]$triangleright visibility
3:$\tilde{C}_{1}\leftarrow C_{1}-\bar{C_{1}}_{S_{t}}$
4:$\tilde{C}_{0}\leftarrow\textsc{ZeroTensor}()$
@@
This process results in data pairings where both $C_{0}$ and $C_{t}$ are properly centered according to atoms that are possibly valid at time $t$. It is important to note that under this scheme, while the model is $SE(3)$-equivariant with respect to the system defined by the partial mask $S_{t}$, it is not equivariant with respect to the orientation of the molecule itself unless $\mathcal{D}_{t}^{c}\;=\;\varnothing$, as the presence and temporary validity of masked dummy atoms offsets the true atom centering and thus breaks both translational and rotational equivariance.
Flexible Atom Count.
Most 3D molecule generation methods require specifying the number of atoms during inference. Because the prior of SynCoGen is over building blocks, we naturally handle a flexible number of atoms during generation and model any excessive atoms as ghost atoms.
4.3 Training-time Constraints
SynCoGen inherits training-time simplifications from MDLM (Sahoo et al., 2024), including zero masked logit probabilities and carry-over logit unmasking. In addition, we implement the following:
  1. 1.
  2. 2.
    Edge Count Limit. Let
    $k_{t}:=\sum_{1\leq i<j\leq n}\mathds{1}\bigl{[}E_{t}[i,j,\cdot]\notin\{\pi_{E},\lambda_{E}\}\bigr{]},$ (4)
    be the number of unmasked true edges in the upper triangle of $E_{t}$. If $k_{t}=n-1$, we have the correct number of edges for a molecule containing $n$ building blocks and therefore set all remaining edge logits to $\lambda_{E}$.
  3. 3.
    Compatibility Masking. Assume that for some $E_{t}$ an edge entry is already denoised, $E_{t}[i,j,\cdot]=(r,v_{i},v_{j})$, meaning that building block $i$ reacts with building block $j$ via reaction $r$ and centers $v_{i}\in\mathcal{V}(X_{i})$, $v_{j}\in\mathcal{V}(X_{j})$. Define the sets of center-matched reagents
    $\textstyle\mathcal{B}_{r,v}^{A}$ :=$:=\{\,b\in\mathcal{B}\mid(b,v)\text{ matches reagent A in }r\},$ (5)
    $\mathcal{B}_{r,v}^{B}$ :=$:=\{\,b\in\mathcal{B}\mid(b,v)\text{ matches reagent B in }r\}.$
    For every node slot $i$ (resp. $j$) we construct a $|\mathcal{B}|$-dimensional binary mask
    $\textstyle\mathcal{X}_{i,k}=\mathds{1}[b_{k}\in\mathcal{B}_{r,v_{i}}^{A}],$ (6)
    $\mathcal{X}_{j,k}=\mathds{1}[b_{k}\in\mathcal{B}_{r,v_{j}}^{B}],$
    $k=1,\dots,|\mathcal{B}|.$
    so that the soft-max for $X_{t}[i,\cdot]$ (resp. $X_{t}[j,\cdot]$) is evaluated only over the 1-entries of $\mathcal{X}_{i}$ (resp. $\mathcal{X}_{j}$). Analogously, once a node identity $X_{t}[j]=b$ is denoised, incoming edge channels $(i,j)$ with $j>i$ are masked to reactions $e=(r,v_{i},v_{j})$ such that $b\in\mathcal{B}_{r,v_{i}}^{B}$.
Put simply, we restrict logits to disallow loops (e.g. macrocycles), to impose a limit on the number of edges, and to better ensure the selection of chemically compatible building blocks and reactions.
4.4 Sampling
Sampling begins by drawing a building block count $n\sim\operatorname{Cat}(\pi_{\text{frag}})$, setting the node and edge tensors to the masked tokens, $X_{1}[i,\!\cdot]=\pi_{X},\;E_{1}[i,j,\!\cdot]=\pi_{E}$ for every $0\leq i,j<N$, and padding all $(i\geq n)$ rows/columns with the no–edge token $\lambda_{E}$. The initial coordinates are an isotropic Gaussian $C_{1}\sim\mathcal{N}(0,I)^{N\times M\times 3}$. From this state, the sampler walks backwards in diffusion time, and at each step it (i) recenters the current coordinates by the visibility mask $S_{t}$ derived from $X_{t}$, (ii) generates node, edge and coordinate predictions with the trained model, (iii) applies the single-parent edge constraint (Algorithm 2), (iv) draws the next discrete state, and (v) updates coordinates by an Euler step. After a final, deterministic pass, we calculate $\bigl{(}\hat{X}_{0},\hat{E}_{0}\bigr{)}=\arg\max_{k}L_{\theta}^{E}[\cdots,k]$ and center the coordinates to yield the molecule $(\hat{X}_{0},\hat{E}_{0},\hat{C}_{0})$. Complete pseudocode is provided in Section B.4. We note our discrete and continuous schemes share a unified time. Lastly, we find inference annealing on the coordinates (see Section D.1) yields small performance gains at sampling time.
Constraint-Aware Edge Pruning.
@@ 4.4 Sampling
By construction, a molecule containing $n$ connected building blocks contains exactly $n-1$ edges, and fragment $j>0$ has a unique parent $i<j$. Consequently, let $E^{t}_{\theta}\in[0,1]^{n\times n\times|\mathcal{R}|V_{\max}^{2}}$ be the soft-max edge probabilities produced at step $t$. The routine below resolves the unique parent for every building block column $j>0$ and returns a probability tensor $\tilde{E}^{t}_{\theta}$ with exactly one non–zero entry per column.
Algorithm 2 SampleEdges$\bigl{(}E^{t}_{\theta},n\bigr{)}$
@@
$\tilde{E}^{t}_{\theta}$ is then symmetrized and fed to the discrete reverse sampler described in Section B.6.
5 Experiments
5.1 De Novo 3D Molecule Generation
For each model, we sample 1000 molecules and compute stringent metrics capturing chemical soundness, synthetic accessibility, conformer quality, and distributional fidelity. Pertaining to the molecular graphs, we report the RDKit sanitization validity (Valid.) and retrosynthetic solve rate (AiZynthFinder (Genheden et al., 2020) (AiZyn.) and Syntheseus (Maziarz et al., 2025) (Synth.)). For conformers, we introduce two physics-based metrics: the median non-bonded interaction energies per atom via the forcefield method GFN-FF and via the semiempirical quantum chemistry method GFN2-xTB Bannwarth et al. (2019); Spicher and Grimme (2020); we also check PoseBusters (Buttenschoen et al., 2024) validity rate (PB). We evaluate the diversity (Div.) as the average pairwise Tanimoto similarity, novelty (Nov.) as the percentage of candidates not appearing in the training set, and the Fréchet ChemNet Distance (Preuer et al., 2018) (FCD) on the distance between generated samples and the training distribution. Details on how the metrics are computed can be found in Section D.2.
Table 1: Comparison of generative methods for de novo 3D molecule generation. ↑/↓ indicate that higher/lower is better.
@@ 5.1 De Novo 3D Molecule Generation
Table 2: Ablation studies. We incrementally remove inference annealing, auxiliary losses, self-conditioning, scaled-noise, and constraints to see the performance difference. All results shown are at 50 epochs rather than 100 epochs in Table 1. See Sections 4.4 and 4.3, (Sections B.9, B.3 and B.11).
Method Valid. $\uparrow$ GFN-FF $\downarrow$
@@
The results are presented in Table 1, and example molecules can be seen in Figure 8. For chemical reasonableness, we see that SynCoGen generates almost entirely valid molecules. Our generation process mimics a multi-step reaction pathway, and as a result, our molecules are significantly more synthesizable compared to baseline methods. Because AiZynthFinder and Syntheseus solve only 50–70 % of known drug-like molecules, our 50–72 % scores likely underestimate true synthesizability.
For structural reasonableness, the generated conformers reproduce the overall energy distributions and have very favorable non-covalent interaction energies as evaluated by relatively accurate computational chemistry methods, especially when compared to the baseline methods (Tables 1 and 3). This is also evident from the lack of structural changes upon further geometric relaxation (Figure 9). The low non-bonded energies indicate SynCoGen learns to sample many intramolecular interactions (Figure 8). Quantitatively, 87% of these conformers pass PoseBusters pose plausibility checks. Furthermore, SynCoGen reproduces the delicate data distribution of bond lengths, angles, and dihedrals (Figures 3 and 7). For example, SynCoGen generates fewer $sp^{2}$-$sp^{2}$ bonds that are too short, captures sharp bond angle distributions (e.g., $sp^{3}$-$sp^{3}$-$sp^{3}$), and replicates the broad dihedral angle distribution of the flexible $sp^{3}$-$sp^{3}$-$sp^{3}$-$sp^{3}$ while still covering the rigid dihedral angles (e.g., $sp^{3}$-$sp^{2}$-$sp^{2}$-$sp^{2}$). Wasserstein-1 distances and JSD can be found in Table 4, and additional plots on other bond lengths/angles/dihedrals can be found in Figure 7.
Figure 3: Conformer geometry and energy distribution comparisons. Distributions of a) bond lengths, b) bond angles, c-d) dihedral angles, e) average per-atom GFN2-xTB non-covalent interaction energies and f) average per-atom GFN-FF non-bonded interaction energies. Solid curves denote training data densities; lower subpanels in (a-d) show deviations between generated samples and data.
Training-, sampling-time and architectural ablations are presented in Tables 2 and D.1. The largest performance enhancement originates from chemistry-aware graph constraints and self-conditioning, with small contributions from other training/sampling details. The performance gap between SynCoGen and SemlaFlow retrained on SynSpace underscores that our training procedure — rather than the architecture or dataset — is the primary driver of performance.
The multi-modal model can further perform other tasks; for example, given randomly assembled molecules from reaction graphs, SynCoGen can perform zero-shot conformer generation at a quality similar to ETKDG as implemented in RDKit (Table 5).
Finally, SynCoGen captures the training distribution as indicated by the low FCD, while generally producing novel molecules. The generated samples have slightly lower diversity as a trade-off for using a (limited) set of reaction building blocks. All generated samples are unique.
5.2 Molecular Inpainting for Fragment Linking
Figure 4: Molecular inpainting. a) Fragment linking starts from three experimentally identified ligands in the PDB that contain substructure matches with our building blocks. We show three examples of linkers generated by SynCoGen per structure and the distribution of Vina docking scores. b) Proposed synthesis pathway for molecule (1) sampled from our model and c) structure of (1) (blue) docked onto PDB 7N7X using AlphaFold3 compared against the PDB ligand (beige).
To demonstrate the practical usefulness of SynCoGen, we study fragment linking (Bancet et al., 2020) to design easily synthesizable analogs of hard-to-make drugs. Fragment linking in drug design enables the construction of potent molecules, such as PROTAC bivalent protein degraders Békés et al. (2022); Chirnomas et al. (2023), by connecting smaller fragments that are known to bind distinct regions of a target site. We formulate fragment linking as a molecular inpainting task, where we fix the identity and coordinates of two fragments in a known ligand and sample its missing parts consistent with both geometry and reaction grammar.
As a case study, we pick several FDA-approved, hard-to-synthesize small molecules with experimental crystal structures, each bound to a different target protein. We select human plasma kallikrein (PDB: 7N7X), multidrug-resistant HIV protease 1 (PDB: 4EYR), and human cyclin-dependent kinase 6 (PDB: 5L2S), where each structure is complexed with a ligand that contains at least two of our building blocks. At sampling time, we condition on the substructure match by keeping fixed fragments denoised and replacing their coordinates with the corresponding linearly interpolated configuration at each time step $t$ (see Section B.13 for details).
We evaluate our generated molecules using AutoDock Vina (Figure 4(Eberhardt et al., 2021). SynCoGen consistently produces molecules with docking scores on par with or better than the native ligand while satisfying constraints on the presence of specific building blocks. AlphaFold3 (Abramson et al., 2024b) folding predictions on selected protein-ligand pairs show similar binding positions in the selected pockets as well (Figures 4 and 10). Crucially, unlike existing approaches (Schneuing et al., 2024; Igashov et al., 2024), the model links fragments using building blocks and reactions to ensure streamlined synthetic routes of the designs (Table 6). Sample multi-step synthetic pathways are shown in Figure 11.
We emphasize that using SynCoGen for fragment-linking does not require retraining. We benchmarked SynCoGen against the state-of-the-art, purpose-built fragment-linking model DiffLinker (Igashov et al., 2024). While SynCoGen produces fewer valid molecules, the valid ones have similar PoseBuster validity rate, lower intramolecular interaction energies, and no disconnected fragments. Most importantly, the retrosynthesis solve rate for DiffLinker is 0% whereas for SynCoGen, it is 58-79% (Table 6). This proof-of-concept inpainting setup is directly applicable for tasks such as scaffold hopping, synthesizable analog generation, or PROTAC design.
6 Conclusion
In this work, we introduced SynCoGen, a unified generative model that jointly samples building-block reaction graphs and atomic coordinates, thereby designing synthesizable molecules directly in Cartesian space. To enable training, we curated SynSpace, a new dataset of readily synthesizable molecules paired with low-energy 3D conformations. SynCoGen achieves state-of-the-art performance across 3D molecular generation benchmarks, while natively returning a tractable synthetic route for each structure. The practical utility of SynCoGen is demonstrated in molecular inpainting to design easily synthesizable drug analogs with geometry constraints.
With SynCoGen, several future directions in synthesizability-aware 3D generation are possible. Beyond chemical graphs and conformers, property-conditioned generation—either through multimodal guidance or direct conditioning on binding pockets—could potentially yield new, high-affinity ligands. Additionally, the current constrained set of reactions and building blocks deliberately favor reliable, automated synthesis at the expense of breadth, and its expansion would open larger regions of chemical space. Beyond drug design, the same co-generative principle may be extended to materials discovery, where synthesizability and 3D accuracy are equally pivotal.
Acknowledgments and Disclosure of Funding
Appendix A Chemistry Details
A.1 Building Blocks and Reactions

### v2

@@ B.6 Data Pairing
Output: $\tilde{C}_{0}$ (re‑centered ground truth), $C_{t}$ (interpolated noisy coords)
[…]
@@
Here, $\mathcal{A}_{i}$ is the set of all atom indices $a$ that constitute true atoms in $X_{0}$. Note that $S_{t}=S_{0}$ for all $t$ where $X_{t}$ contains no masked building blocks.
@@
Our data pairings result in both $C_{0}$ and $C_{t}$ that are properly centered according to atoms that are possibly valid at time $t$. It is important to note that under this scheme, while the model is $SE(3)$-equivariant with respect to the system defined by the partial mask $S_{t}$, it is not equivariant with respect to the orientation of the molecule itself unless $\mathcal{D}_{t}^{c}\;=\;\varnothing$, as the presence and temporary validity of masked dummy atoms offsets the true atom centering and thus breaks both translational and rotational equivariance.
B.7 Training Algorithm
Algorithm 3 Training step for SynCoGen
[…]
B.8 Sampling Algorithm
Algorithm 4 Sampling procedure for SynCoGen
[…]
3:for $t=1$ down to 0 do
4:  $\tilde{C}_{t}\!\leftarrow\!C_{t}-\bar{C}_{t,S_{t}}$;
6:  $\tilde{L}^{E}_{t}\!\leftarrow\!\textsc{SampleEdges}(L^{E}_{t},n)$ $\triangleright$ enforce one parent per building block (Algorithm˜5)
7:  $X_{t-\Delta t}\!\leftarrow\!\textsc{CatSample}(L^{X}_{t})$; $E_{t-\Delta t}\!\leftarrow\!\textsc{CatSample}(\tilde{L}^{E}_{t})$ $\triangleright$ take reverse step (Section˜B.11)
10:end for
[…]
13:return $(\hat{X}_{0},\hat{E}_{0},\hat{C}_{0})$
B.9 Inference-Time Edge Constraints
@@
Let $E^{t}_{\theta}\in[0,1]^{n\times n\times|\mathcal{R}|V_{\max}^{2}}$ be the soft‑max edge probabilities produced at step $t$. The routine below resolves the unique parent for every building block column $j>0$ and returns a probability tensor $\tilde{E}^{t}_{\theta}$ with exactly one non–zero entry per column.
[…]
@@
$\tilde{E}^{t}_{\theta}$ is then symmetrized and fed to the discrete reverse sampler described in Section˜B.11.
B.10 Building Block Logit Predictions
The SemlaFlow(Irwin et al., 2025) backbone outputs atom–atom edge features $E^{\mathrm{atom}}_{\theta}\in\mathbb{R}^{B\times(NM)\times(NM)\times d_{\text{edge}}}$. To obtain building block‑level tensors, we apply two parallel 2‑D convolutions (one for nodes, one for edges) with stride $M$, followed by MLP classifiers that map the pooled features back to their original one‑hot vocabularies. Note that the presented model is trained to predict a maximum of 5 building blocks, where the sizes of the molecules (average 566 Da) are near upper limits of molecular weights for typical drug like molecules.
Stride‑pooled convolution.
$\operatorname{Conv2d}(d_{\text{edge}}\!\to d_{\text{edge}},\,k=M,\,s=M)\;\xrightarrow{\text{SiLU}}\;\operatorname{Conv2d}(d_{\text{edge}}\!\to d_{\text{edge}},\,k=1,\,s=1),$ (9)
Node head.
We flatten $X_{\text{pool}}$ along its channel axis, concatenate the residual building block one‑hot matrix $X_{t}$, and pass the result through a two‑layer MLP to obtain
$L_{\theta}^{X_{t}}\in\mathbb{R}^{B\times N\times|\mathcal{B}|}.$ (11)
Edge head.
We concatenate $E_{\text{pool}}$ with the residual building block‑edge one‑hot tensor $E_{t}$, apply an analogous two‑layer MLP, and symmetrize to produce
Atom Features.
The SemlaFlow(Irwin et al., 2025) backbone additionally outputs atom-level node features $X^{\mathrm{atom}}_{\theta}\in\mathbb{R}^{B\times(NM)\times d_{\text{node}}}$, which are incorporated into $E^{\mathrm{atom}}_{\theta}$ via a bond refinement message-passing layer. We find that extracting both building block and edge logits directly from the refined features $E^{\mathrm{atom}}_{\theta}$ marginally improves performance relative to separately predicting $L_{\theta}^{X_{t}}$ from $X^{\mathrm{atom}}_{\theta}$ and $L_{\theta}^{E_{t}}$ from $E^{\mathrm{atom}}_{\theta}$.
B.11 Discrete Noising Scheme
Following (Sahoo et al., 2024), we adopt an absorbing (masked) state noising scheme for $X_{0}$ and $E_{0}$:
where $(\alpha_{t})_{t\in[0,1]}$ is the monotonically decreasing noise schedule introduced in Section˜2.
Reverse categorical posterior.
For node identities, we have
and, analogously, for edge labels
@@ B.11 Discrete Noising Scheme
where $s<t$. Equations˜14 and 15 are the direct translation of the reverse denoising process described by (Sahoo et al., 2024) into SynCoGen’s node–edge representation.
B.12 Noise Schedule Parameterization
Following MDLM (Sahoo et al., 2024), we parameterize the discrete noising schedule via $\alpha_{t}=e^{-\sigma(t)}$, where $\sigma(t):[0,1]\to\mathbb{R}^{+}$. In all experiments, we adopt the linear schedule:
$\sigma(t)=\sigma_{\max}t,$ (16)
Edge Symmetrization.
After drawing the upper‑triangle entries of the one‑hot edge tensor $E_{s}$ in either the forward or reverse (de)noising process, we enforce symmetry by copying them to the lower triangle:
$E_{s,jie}\;=\;E_{s,ije},\qquad 0\leq i<j<n,\;\;e\in\mathcal{R}V_{\max}^{2}.$
B.13 Positional Embeddings
Though SemlaFlow(Irwin et al., 2025) is permutationally invariant by design with respect to atom positions, SynCoGen dataset molecules require that atom order be fixed and grouped by building block for reconstruction purposes. To enforce this during training, we intentionally break permutation invariance by generating and concatenating to each input coordinate sinusoidal positional embeddings representing both global atom index and building block index.
B.14 Hyperparameters
We train SynCoGen for 100 epochs with a batch size of 128 and a global batch size of 512. Note that SemlaFlow and Midi are trained for 200 epochs, and EQGAT-diff is trained for up to 800 epochs. All models are trained with a linear noise schedule (see Section˜B.12), with the SUBS parameterization enabled. During training, a random conformer for each molecule is selected, then centered and randomly rotated to serve as the ground-truth coordinates $C_{0}$. All atomic coordinates are normalized by a constant $Z_{c}$ describing the standard deviation across all training examples. For the pairwise distance loss $\mathcal{L}_{pair}$, we set $d$ to 3Å, adjusted for normalization. During training, for each recentered input-prior pair $(\tilde{C}_{1},\tilde{C}_{0})$ we rotationally align $C_{1}$ to $C_{0}$. When training with noise scaling and the bond loss time threshold, we set the noise scaling coefficient to 0.2 and the time threshold to 0.25, above which bond length losses are zeroed. When training with auxiliary losses, we set the weights for the pairwise, sLDDT, and bond length loss components to 0.4, 0.4, and 0.2, respectively.
B.15 Computational Resources Used
We train all models on 2 H100-80GB GPUs.
B.16 Training Losses
Here, we define several loss terms that have proved useful for stabilizing training on 3‑D geometry.
Graph loss.
Let $X_{0}$ and $E_{0}$ be the clean node and edge tensors. Following the MDLM implementation (Sahoo et al., 2024), we weigh the negative log‑likelihood at step $t$ by
i.e. the cross‑entropy between the one‑hot ground truth and the predicted distributions for both nodes and edges.
MSE loss.
Pairwise loss.
where $d$ is the distance cut‑off for pairwise terms. The default total loss value for the model is therefore
Smooth‑LDDT loss (Abramson et al., 2024a).
Let $d_{ij}^{0}:=\lVert C_{0}[i]-C_{0}[j]\rVert_{2}$ and $d_{ij}^{\text{pred}}:=\lVert\hat{C}_{0}[i]-\hat{C}_{0}[j]\rVert_{2}$ be ground‑truth and predicted inter‑atomic distances, respectively. For each pair of atoms within a $15$ Å cutoff in the reference structure, we compute the per‑pair score
where $\sigma(x)=1/(1+e^{-x})$ is the logistic function. The smooth‑LDDT loss averages $1-\operatorname{sLDDT}_{ij}$ over all valid pairs,
Bond‑length loss.
Given a set of intra‑fragment bonds $\mathrm{bonds}=\{(p,q)\}$ extracted from the vocabulary, we penalize deviations in predicted bond lengths:
Self-Conditioning.
[…]
B.17 Conformer generation
We randomly assembled 50 molecules with the reaction graph and used the standard conformational search (iMTD-GC) in CREST with GFN-FF to find all reference conformers. For both SynCoGen and RDKit ETKDG, we sampled 50 conformers per molecule and computed the coverage and matching scores. We used a relatively strict RMSD threshold of $\tau=0.75\ \text{\AA }$.
Formally, COV is defined as:
B.18 Molecular inpainting
For the inpainting experiments in Section˜5.2, we keep two fragments $\mathcal{D}=\{\mathcal{D}^{(1)},\mathcal{D}^{(2)}\}$ and their coordinates fixed and sample the remaining part of the molecule. We follow Section˜B.8 and initialize the graph prior $X_{1}$ with the one-hot encoding of the desired fragment $i$ at a specified node index in the graph (decided at random or based on the structure of the original molecule, so that it matches its scaffold). For each denoised fragment $\mathcal{D}^{(i)}$, we replace its coordinates at each time $t>0.03$ during sampling by
$C_{t}^{(i)}\;=\;(1-t)\,\tilde{C}_{0}^{(i)}\;+\;t\,\tilde{C}_{1}^{(i)},$
where $\tilde{C}_{0}^{(i)}$ and $\tilde{C}_{1}^{(i)}$ are the centered ground-truth and prior coordinates of fragment $i$, respectively, and all other fragments are updated as shown in Section˜B.8. For any $t\leq 0.03$, which for 100 sampling steps amounts to the last three steps in the path, we follow normal Euler steps as shown in Section˜B.8 to allow a refinement of the fixed coordinates in line with the rest of the predicted ones for the rest of the fragments. We empirically observed that this led to molecules with lower average energies.
Appendix C Baseline comparisons.
C.1 Unconditional Generation.
SemlaFlow
We evaluated SemlaFlow using the sampling script in the official codebase on GitHub [^1] https://github.com/rssrwn/semla-flow/, available under the MIT License. We reported results for a model trained on the GEOM (Axelrod and Gomez-Bombarelli, 2022) dataset (by sampling from the checkpoints provided in the repository) and from a model trained on our dataset (see Table˜1). We trained SemlaFlow using the default hyperparameters for 150 epochs on a single conformer per molecule.
EQGAT-diff, MiDi, JODO, FlowMol
C.2 Conditional Generation.
In the pharmacophore-conditioned generation setting, we compare SynCoGen against Synformer (Jocys et al., 2024), CGFlow (Shen et al., 2025), and ShEPhERD (Adams et al., 2025). SynFormer was conditioned on the native ligand for synthesizable analogue generation. We used the official implementation on GitHub [^3] https://github.com/wenhao-gao/synformer, and changed the following inference settings to allow for higher quality designs compared to the default: search_width=32, exhaustiveness=128, time_limit=300. For ShEPhERD, we use the $p(x_{1}|x_{3},x_{4})$ conditional setting from the paper experiments where $x_{1}$ denotes molecular structure, $x_{3}$ denotes the reference ligand charge surface, and $x_{4}$ denotes the reference ligand pharmacophore profile. We provide ShEPhERD with the reference ligand and generate 100 analogs evenly split between 36, 38, 40, 42, 44, 46, 48, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 62, 64, 66, 68, 70, 75, and 80 atoms (4 each). To prepare molecules for ShEPhERD conditioning, we generate partial charges for each reference ligand using xTB(Bannwarth et al., 2019). For CGFlow-ZS experiments, we generate molecules in a zero-shot protein-conditioned setting using TacoGFN (Shen et al., 2024) first implemented by Seo et al.(Seo et al., 2024; Shen et al., 2024) Molecules are generated using the web app described in the GitHub repository [^4] https://github.com/tsa87/cgflow, which inherits the sampling hyperparameters specified in the original CGFlow manuscript. For each target, we conditionally generate using a cleaned PDB and centroid derived from reference ligand heavy atoms. See  Section˜D.6 for details of DiffLinker in linker design experiments.
Appendix D Extended results and discussion
D.1 Training Ablations
@@
Table 3: Training ablations. We incrementally remove inference annealing, auxiliary losses, self-conditioning, scaled-noise, and constraints to see the performance difference. All results shown are at 50 epochs rather than 100 epochs in Table˜1. Here, "Constraints" refers to both training-time compatibility masking and sampling constraints. See Sections˜4.4 and 4.3, (Sections˜B.14, B.7 and B.16).
Method Valid. $\uparrow$ GFN-$F$ $\downarrow$
@@
D.2 Sampling Ablations
We additionally follow FoldFlow to implement inference annealing, a time-dependent scaling on Euler step size that was found to empirically improve in-silico results in protein design Bose et al. (2024). We studied multiplying the Euler step size at time $t$ by $5t$, $10t$, and $50t$. In practice, we employ $10t$ for our experiments unless otherwise noted.
Table 4: Sampling ablations. Results are averaged over 1000 generated samples, except retrosynthesis solve rate (out of 100). All results shown are at 50 epochs rather than 100 epochs in Table˜1.